OpenStudy (anonymous):

Express the required measurement in exact form (reduce fractions completely). What part of the circumference is an arc whose measure is a. 1°? c. 45°? e. 90°? b. 30°? d. 60°?

OpenStudy (anonymous):

\[\left\{\left\{1{}^{\circ},\frac{1}{360}\right\},\left\{45 {}^{\circ},\frac{1}{8}\right\},\left\{90 {}^{\circ},\frac{1}{4}\right\},\left\{30 {}^{\circ},\frac{1}{12}\right\},\left\{60 {}^{\circ},\frac{1}{6}\right\}\right\} \]

OpenStudy (anonymous):

do i need to divide them like 1/360 or is that the answeer

OpenStudy (anonymous):

No division is required. They are in the form as requested in the problem statement.

OpenStudy (anonymous):

The fractions are the result of dividing each of your angles by 360 degrees respectively.